# DPLib: A Standard Benchmark Library for Distributed Power System Analysis and Optimization

### v2

@@ III-A Distributed OPF Formulation
We formulate the OPF problem in both DC and AC settings for each region $r$. Each region solves a local OPF problem while maintaining consistency at the boundaries via duplicated (virtual) variables and tie-line constraints. These virtual variables (buses) enable each region to model tie-line interactions locally without requiring direct enforcement of global constraints. The overall structure is illustrated in Fig. 1(a), where a two-region power grid is shown with inter-regional tie-lines in red. Fig. 1(b) further illustrates how virtual buses (shown in red) are added to represent shared variables.
@@ III-A1 DC OPF Formulation
$\qquad\quad\underline{p}_{ij}\leqslant p_{ij}\leqslant\bar{p}_{ij},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (1d)
$\qquad\quad\theta_{\text{ref }}=0,$ (1e)
$\qquad\quad\theta_{i}=\tilde{\theta}_{i},\quad\tilde{\theta}_{j}=\theta_{j},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (1f)
$\boldsymbol{x}^{r}=[\boldsymbol{p}^{g\top},\boldsymbol{\theta}^{\top}]\in\mathbb{R}^{n^{r}_{g}+n^{r}_{b}}$ is the decision vector for region $r$, including generation $p^{g}_{i}$ and voltage angles $\theta_{i}$. Equation (1a) minimizes the local cost $f(p^{g}_{i})$. Constraints (1b)–(1c) enforce generator limits and power balance using demand $p^{d}_{i}$, line reactance $x_{ij}$, and virtual angles $\tilde{\theta}_{j}$. Constraint (1d) bounds internal and tie-line flows $p_{ij}$, and (1f) ensures consistency between real and virtual bus angles.
@@ III-A2 AC OPF Formulation
$\qquad\underline{p}_{i}^{g}\leqslant p^{g}_{i}\leqslant\bar{p}_{i}^{g},\quad\underline{q}_{i}^{g}\leqslant q^{g}_{i}\leqslant\bar{q}_{i}^{g},\quad\forall i\in\Omega^{r}_{g}$ (2b)
$\qquad p^{g}_{i}-p^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\left(p_{ij,f}+p_{ij,t}\right)+\sum_{j\in\Omega^{r,\text{in}}_{i}}\tilde{p}_{ij},\quad\forall i\in\Omega^{r}_{b}$ (2c)
$\qquad q^{g}_{i}-q^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\left(q_{ij,f}+q_{ij,t}\right)+\sum_{j\in\Omega^{r,\text{in}}_{i}}\tilde{q}_{ij},\quad\forall i\in\Omega^{r}_{b}$ (2d)
@@
$\qquad\qquad\quad+B_{ij,2}v_{i}v_{j}\sin(\theta_{i}-\theta_{j}),\quad\forall ij\in\Omega^{r}_{l}$ (2e)
@@
$\qquad\qquad\quad+G_{ij,2}v_{i}v_{j}\sin(\theta_{i}-\theta_{j}),\quad\forall ij\in\Omega^{r}_{l}$ (2f)
@@
$\qquad\qquad\quad+B_{ij,3}v_{i}v_{j}\sin(\theta_{j}-\theta_{i}),\quad\forall ij\in\Omega^{r}_{l}$ (2g)
@@
$\qquad\qquad\quad+G_{ij,3}v_{i}v_{j}\sin(\theta_{j}-\theta_{i}),\quad\forall ij\in\Omega^{r}_{l}$ (2h)
$\qquad p_{ij,f}^{2}+q_{ij,f}^{2}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (2i)
$\qquad p_{ij,t}^{2}+q_{ij,t}^{2}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (2j)
@@
$\qquad\underline{\Delta\theta}_{ij}\leq\theta_{i}-\theta_{j}\leq\overline{\Delta\theta}_{ij},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (2k)
$\qquad\theta_{\text{ref}}=0$ (2l)
@@
$\qquad\theta_{i}=\tilde{\theta}_{i},\quad\tilde{\theta}_{j}=\theta_{j},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (2m)
$\qquad v_{i}=\tilde{v}_{i},\quad\tilde{v}_{j}=v_{j},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (2n)
The decision vector is $\boldsymbol{x}^{r}=[\boldsymbol{p}^{g\top},\boldsymbol{q}^{g\top},\boldsymbol{\theta}^{\top},\boldsymbol{v}^{\top}]$, including generation $(p^{g}_{i},q^{g}_{i})$, angles $\theta_{i}$, and voltages $v_{i}$ at region $r$. Constraints (2b)–(2d) enforce generator and nodal power limits; (2e)–(2h) define AC flows using admittance terms $G_{ij,k},B_{ij,k}$. Limits on flows and angle differences are imposed in (2i)–(2k). Reference angle (2l) and virtual bus equalities (2m)–(2n) ensure boundary consistency.
III-A3 ADMM Coordination
To ensure consistency between neighboring regions, we use ADMM to coordinate virtual variables [10]. For example, the voltage angle consistency is enforced via:
$L_{\rho}(\theta,\tilde{\theta},\lambda)=f(p^{g})+\lambda^{\top}(\theta-\tilde{\theta})+\frac{\rho}{2}\|\theta-\tilde{\theta}\|^{2},$ (3)
where $\lambda$ are the dual variables and $\rho$ is the penalty parameter. Similar terms are added for voltage magnitudes in AC OPF.
Each ADMM iteration proceeds as:
  1. 1.
    Solve each region’s local OPF using the current dual variables.
  2. 2.
    Update dual variables based on tie-line consistency.
  3. 3.
    Check primal residuals and stop if below a threshold.
In general, coordination across regions is achieved using the ADMM. The distributed DC OPF and AC OPF formulations in (1) and (2) incorporate local objectives and tie-line consistency via augmented Lagrangian. Subproblems are solved locally at each iteration, dual variables are updated, and convergence is checked via the worst primal residual.
@@ III-B1 DPLib DC OPF Solver
The DPLib DC OPF solves the OPF problem using ADMM by optimizing each region independently and enforcing agreement on shared variables (tie-line voltage angles) through iterative dual updates. The solver is available at GitHub, including the MATLAB implementation file named DistDCOPF.m. Designed for partitioned MATPOWER cases, it requires MATLAB R2020a or newer with MATPOWER and YALMIP installed. Upon execution, the user provides the partitioned .mat file, ADMM penalty parameter $\rho$, convergence threshold, maximum number of iterations, and optionally the centralized OPF cost for benchmarking via optimality gap.
After receiving user inputs, the code loads all regional and tie-line data, augments each region’s branch matrix with inter-regional tie-lines, and initializes bus voltage angle variables and indexing. The core ADMM loop then solves each region’s DC OPF using YALMIP, minimizing generation cost under local constraints (e.g., power balance, generator limits) while adding consistency penalties on shared tie-line angles as modeled in (1). A helper function handles the regional problem formulation and solution.
After each iteration, the code evaluates tie-line angle consistency and updates dual variables accordingly. Primal residuals are recorded, and convergence is declared if the maximum error drops below the user-defined threshold; otherwise, the process continues until the iteration limit. The optimality gap is computed and plotted alongside residuals if a centralized cost is provided. Progress and convergence status are printed in the command window.
The DPLib DC OPF solver supports any number of regions and is tailored for use with partitioned .mat files, enabling reproducible distributed OPF testing. Algorithm 2 summarizes its workflow.
Algorithm 2 DPLib OPF solver
1:Input: Partitioned MATPOWER data, ADMM parameters $\rho$, convergence tolerance $\epsilon$, maximum iterations
2: Load the partitioned case data for all regions
3: Augment each region with its corresponding tie-line data
4: Initialize region variables including local decision variables and dual variables
5: Retrieve ADMM parameters and consistency constraints
6: Identify and prepare interface buses at region boundaries
7:for each ADMM iteration do
8:  Solve local OPF for each region independently
9:  Update dual variables based on consistency violations at bordering buses
10:  if convergence criteria are met then
11:   break
12:  end if
13:end for
14: Visualize and save final solution

### v3

@@ III-A Distributed OPF Formulation
We formulate the OPF problem in both DC and AC settings for each region $r$. Each region solves a local OPF problem while maintaining consistency at the boundaries via duplicated (virtual) variables and tie-line constraints [8]. These virtual variables (buses) enable each region to model tie-line interactions locally without requiring direct enforcement of global constraints. The overall structure is illustrated in Fig. 1(a), where a two-region power grid is shown with inter-regional tie-lines in red. Fig. 1(b) further illustrates how virtual buses (shown in red) are added to represent shared variables.
@@ III-A1 DC OPF Formulation
$\qquad\quad\underline{p}_{ij}\leqslant\frac{\theta_{i}-\theta_{j}}{x_{ij}}\leqslant\bar{p}_{ij},\quad\forall ij\in\Omega^{r}_{l}$ (3d)
$\qquad\quad\underline{p}_{ij}\leqslant\frac{\theta_{i}-\tilde{\theta}_{j}}{x_{ij}}\leqslant\bar{p}_{ij},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (3e)
$\qquad\quad\theta_{\text{ref }}=0,$ (3f)
$\qquad\quad\tilde{\theta}_{j}=\theta_{j},\quad\forall j:ij\in\Omega^{r,\textrm{in}}_{l}$ (3g)
$\qquad\quad\theta_{i}=\tilde{\theta}_{i},\quad\forall i:ij\in\Omega^{r,\textrm{in}}_{l}$ (3h)
The decision vector for region r contains the active power generations $p^{g}_{i}$ for all generating units $i\in\Omega^{r}_{g}$ and the bus voltage phase angles $\theta_{i}$ for all buses $i\in\Omega^{r}_{b}$, giving $x^{r}\in\mathbb{R}^{n^{r}_{g}+n^{r}_{b}}$. The objective in (3a) minimizes the local generation cost $f(p^{g}_{i})$. Constraint (3b) applies the minimum and maximum generation limits for all units in the set $\Omega^{r}_{g}$. The nodal balance equation (3c) uses the demand $p^{d}_{i}$, the reactance $x_{ij}$ of each line, the set of internal neighboring buses $\Omega^{r}_{i}$, and the set of inter-regional neighboring buses $\Omega^{r,\mathrm{in}}_{i}$, and incorporates virtual boundary angles $\tilde{\theta}_{j}$ that represent buses belonging to adjacent regions. Constraints (3d)–(3e) impose the power-flow limits $\underline{p}_{ij}$ and $\bar{p}_{ij}$ on all internal lines in $\Omega^{r}_{l}$ and all tie-lines in $\Omega^{r,\mathrm{in}}_{l}$. Constraint (3f) selects the regional reference bus. Finally, constraints (3g)–(3h) enforce angle consistency across inter-regional connections by matching the real boundary-bus angles with their corresponding virtual angles.
@@ III-A2 AC OPF Formulation
$\qquad\underline{p}_{i}^{g}\leqslant p^{g}_{i}\leqslant\bar{p}_{i}^{g},\quad\forall i\in\Omega^{r}_{g}$ (4b)
$\qquad\underline{q}_{i}^{g}\leqslant q^{g}_{i}\leqslant\bar{q}_{i}^{g},\quad\forall i\in\Omega^{r}_{g}$ (4c)
$\qquad p^{g}_{i}-p^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\left(p_{ij,f}+p_{ij,t}\right)+\sum_{j\in\Omega^{r,\text{in}}_{i}}{p}_{ij,\xi},\quad\forall i\in\Omega^{r}_{b}$ (4d)
$\qquad q^{g}_{i}-q^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\left(q_{ij,f}+q_{ij,t}\right)+\sum_{j\in\Omega^{r,\text{in}}_{i}}{q}_{ij,\xi},\quad\forall i\in\Omega^{r}_{b}$ (4e)
@@
$\qquad\qquad\quad+B_{ij,2}v_{i}v_{j}\sin(\theta_{i}-\theta_{j}),\quad\forall ij\in\Omega^{r}_{l}$ (4f)
@@
$\qquad\qquad\quad+G_{ij,2}v_{i}v_{j}\sin(\theta_{i}-\theta_{j}),\quad\forall ij\in\Omega^{r}_{l}$ (4g)
@@
$\qquad\qquad\quad+B_{ij,3}v_{i}v_{j}\sin(\theta_{j}-\theta_{i}),\quad\forall ij\in\Omega^{r}_{l}$ (4h)
@@
$\qquad\qquad\quad+G_{ij,3}v_{i}v_{j}\sin(\theta_{j}-\theta_{i}),\quad\forall ij\in\Omega^{r}_{l}$ (4i)
$\qquad p_{ij,f}^{2}+q_{ij,f}^{2}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r}_{l}$ (4j)
$\qquad p_{ij,t}^{2}+q_{ij,t}^{2}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r}_{l}$ (4k)
$\qquad{p}^{2}_{ij,\xi}+{q}^{2}_{ij,\xi}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (4l)
@@
$\qquad\underline{\Delta\theta}_{ij}\leq\theta_{i}-\theta_{j}\leq\overline{\Delta\theta}_{ij},\quad\forall ij\in\Omega^{r}_{l}$ (4m)
@@
$\qquad\underline{\Delta\theta}_{ij}\leq\theta_{i}-\tilde{\theta}_{j}\leq\overline{\Delta\theta}_{ij},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (4n)
$\qquad\theta_{\text{ref}}=0,$ (4o)
$\qquad\tilde{\theta}_{j}=\theta_{j},\quad\forall j:ij\in\Omega^{r,\textrm{in}}_{l}$ (4p)
$\qquad\theta_{i}=\tilde{\theta}_{i},\quad\forall i:ij\in\Omega^{r,\textrm{in}}_{l}$ (4q)
$\qquad\tilde{v}_{j}=v_{j},\quad\forall j:ij\in\Omega^{r,\textrm{in}}_{l}$ (4r)
$\qquad v_{i}=\tilde{v}_{i},\quad\forall i:ij\in\Omega^{r,\textrm{in}}_{l}$ (4s)
The AC OPF subproblem for region r uses the same generation set $\Omega^{r}_{g}$ and bus set $\Omega^{r}_{b}$, where the decision variables include active and reactive generations $p^{g}_{i}$ and $q^{g}_{i}$, bus voltage magnitudes $v_{i}$, and phase angles $\theta_{i}$, together with virtual boundary variables $\tilde{\theta}_{j}$ and $\tilde{v}_{j}$ for all inter-regional neighbors in $\Omega^{r,\mathrm{in}}_{i}$. The objective in (4a) minimizes the generation cost over all units in $\Omega^{r}_{g}$. Generator active and reactive limits are enforced in (4b)–(4c). Power-balance constraints (4d)–(4e) use local demands $p^{d}_{i}$ and $q^{d}_{i}$, as well as flows from internal neighbors $j\in\Omega^{r}_{i}$ and inter-regional neighbors $j\in\Omega^{r,\mathrm{in}}_{i}$, where $p_{ij,f},p_{ij,t},q_{ij,f},q_{ij,t}$ represent the AC branch flows defined by the conductance and susceptance parameters $G_{ij,k}$ and $B_{ij,k}$ in (4f)–(4i). Thermal limits for internal lines $\Omega^{r}_{l}$ and tie-lines $\Omega^{r,\mathrm{in}}_{l}$ are enforced in (4j)–(4l). The phase-angle difference bounds in (4m)–(4n) use the parameters $\underline{\Delta\theta}_{ij}$ and $\overline{\Delta\theta}_{ij}$, and (4o) sets the reference bus angle. Finally, constraints (4p)–(4s) ensure consistency between the real and virtual boundary-bus variables, matching $\theta_{i}$ with $\tilde{\theta}_{i}$ and $v_{i}$ with $\tilde{v}_{i}$ across all inter-regional connections in $\Omega^{r,\mathrm{in}}_{l}$.
III-A3 ADMM Decomposition
The distributed OPF solver uses the sequential form of ADMM, where regions are updated one after another in a Gauss–Seidel manner. Each region solves its local OPF problem using the most recent boundary voltage information from neighbors that have already been updated in the current iteration, while using the previous-iteration values for neighbors that have not yet been updated. The shared variables are the voltage phase angles and voltage magnitudes of boundary buses. Their virtual copies represent the neighboring regions’ views of these boundary variables.
Before the ADMM iterations start, we also compute problem-dependent scaling factors for the boundary mismatches. For the phase angles, a characteristic scale $\theta_{\text{scale}}$ is obtained from the line reactances and thermal limits as
$\theta_{\max,\ell}\approx|x_{\ell}|\,\frac{\text{rateA}_{\ell}}{\text{baseMVA}},$
and $\theta_{\text{scale}}$ is chosen as a robust statistic (for example, the median) of $\{\theta_{\max,\ell}\}$ over all lines. For AC OPF, an additional voltage-magnitude scale $v_{\text{scale}}$ is defined from the bus voltage limits, for example as half the spread between the minimum and maximum allowed magnitudes. These scales are used to normalize the consensus residuals and make them dimensionless.
Consistency across regions is enforced by an augmented Lagrangian of the form
$L_{\rho}(\theta,\tilde{\theta},v,\tilde{v},\lambda_{\theta},\lambda_{v})=$ $f(p^{g})+\lambda_{\theta}^{\top}(\frac{\theta-\tilde{\theta}}{\theta_{\text{scale}}})+\lambda_{v}^{\top}(\frac{v-\tilde{v}}{v_{\text{scale}}})$
$+\tfrac{\rho_{\theta}}{2}\|\frac{\theta-\tilde{\theta}}{\theta_{\text{scale}}}\|^{2}+\tfrac{\rho_{v}}{2}\|\frac{v-\tilde{v}}{v_{\text{scale}}}\|^{2},$ (5)
[…]
During iteration $k$, region $r$ performs the primal update by solving its local AC or DC OPF subproblem. The virtual variables from neighbors are fixed parameters in this optimization. After the OPF solve, region $r$ updates its real boundary voltages $(\theta_{r}^{k+1},v_{r}^{k+1})$ and refreshes its virtual copies $(\tilde{\theta}_{r}^{k+1},\tilde{v}_{r}^{k+1})$ using the newest available values from neighboring regions. This constitutes the primal update step of ADMM. Once all regions finish their primal updates, the dual variables are updated based on normalized residuals. For each tie-line boundary bus, the updates are
$\lambda_{\theta}^{k+1}$ $=\lambda_{\theta}^{k}+\rho_{\theta}\frac{\theta^{k+1}-\tilde{\theta}^{k+1}}{\theta_{\text{scale}}},$ (6)
$\lambda_{v}^{k+1}$ $=\lambda_{v}^{k}+\rho_{v}\frac{v^{k+1}-\tilde{v}^{k+1}}{v_{\text{scale}}}.$ (7)
In the DC-OPF case only the angle channel is present, so the update for $\lambda_{v}$ and the associated terms are omitted.
Convergence is assessed using the worst-case normalized primal residual across all tie-lines. The primal residual at iteration $k+1$ is computed as
$r_{\text{pri}}^{k+1}=\max_{(i,j)\in\mathcal{E}_{\text{tie}}}\left\{\frac{|\theta_{i}^{k+1}-\tilde{\theta}_{i}^{k+1}|}{\theta_{\text{scale}}},\,\frac{|v_{i}^{k+1}-\tilde{v}_{i}^{k+1}|}{v_{\text{scale}}}\right\},$
matching the residual definition used in [8], up to the normalization by $(\theta_{\text{scale}},V_{\text{scale}})$. In the DC-OPF case only the angle term is used in the maximum. The ADMM iterations terminate once this residual falls below a prescribed tolerance.
To improve robustness across different network topologies and operating points, the penalty parameters are updated adaptively based on the relative sizes of the primal and dual residuals, following the heuristic in [17]. Let $r_{\text{pri}}^{k}$ and $r_{\text{dual}}^{k}$ denote the normalized primal and dual residuals at iteration $k$. For the distributed DC-OPF solver, a single penalty parameter is updated according to
$\rho^{k+1}=\begin{cases}\tau_{\text{incr}}\,\rho^{k},&\text{if }r_{\text{pri}}^{k}>\mu\,r_{\text{dual}}^{k},\\[1.99997pt] \rho^{k}/\tau_{\text{decr}},&\text{if }r_{\text{dual}}^{k}>\mu\,r_{\text{pri}}^{k},\\[1.99997pt] \rho^{k},&\text{otherwise},\end{cases}$
with chosen parameters $\mu$, $\tau_{\text{incr}}$, and $\tau_{\text{decr}}$. For the distributed AC-OPF solver, the same heuristic is applied separately to the angle and magnitude channels, i.e., to $\rho_{\theta}$ and $\rho_{v}$, using different balance factors for the $\theta$ and $v$ residuals.
To prevent numerical blow-up and avoid instability caused by excessively large penalties, we enforce an upper bound on the adaptive updates. Specifically, the penalty updates are terminated once the penalty reaches a prescribed maximum value: $\rho_{\mathrm{DC}}^{\max}$ for the distributed DC-OPF solver and $\rho_{\mathrm{AC}}^{\max}$ for the distributed AC-OPF solver. In practice, if an update would increase the current penalty beyond the corresponding cap, we set $\rho\leftarrow\rho_{\mathrm{DC}}^{\max}$ (DC) or $\rho_{\theta},\rho_{v}\leftarrow\rho_{\mathrm{AC}}^{\max}$ (AC) and disable further penalty adaptation thereafter. This safeguard ensures the algorithm does not diverge due to runaway growth of $\rho$.
Moreover, to avoid overly frequent parameter changes—which can introduce oscillations and hinder convergence—we apply a cooldown period after each penalty update. Concretely, once $\rho$ (DC) or $\rho_{\theta},\rho_{v}$ (AC) is updated, we postpone any subsequent residual-balance checks and penalty updates for the next $N_{\mathrm{wait}}$ iterations, where $N_{\mathrm{wait}}=20$ for cases with fewer than $2000$ buses and $N_{\mathrm{wait}}=40$ for cases with $2000$ buses or more. This heuristic empirically stabilizes the convergence behavior by allowing the primal and dual iterates to settle under the new penalty before further adaptation is attempted.
Overall, this adaptive strategy automatically increases the penalty when the primal residual dominates and decreases it when the dual residual dominates, promoting a better balance between consensus enforcement and dual convergence while remaining numerically stable through the maximum-penalty cap and the update cooldown mechanism.
Figure 2 summarizes the sequential ADMM workflow, highlighting the order of local OPF solves, voltage exchange, dual updates (with scaled and adaptively tuned penalties), and residual checks.
Initialize ADMMFor $r=1,\dots,K$: solve regional OPF and update boundary voltagesUpdate dual variables on tie-lines (scaled residuals, adaptive $\rho$)Compute worst-case normalized primal residualConverged?StopNext ADMM iterationyesno
Figure 2: Sequential ADMM loop for distributed OPF with scaled residuals and adaptive penalties
@@ III-B1 DPLib DC OPF Solver
The DPLib DC OPF module implements a distributed DC OPF solver based on the ADMM. The MATLAB implementation, available in the file DistDCOPF.m and publicly accessible on GitHub, is designed for partitioned MATPOWER cases and supports the YALMIP optimization backends in MATLAB. The solver requires MATLAB R2020a or newer with MATPOWER installed; YALMIP must also be added to the MATLAB path.
At execution time, the user provides the partitioned MATPOWER file, the ADMM penalty parameter rho, a convergence tolerance, a maximum number of iterations, a $\rho_{\mathrm{DC}}^{\max}$, and optionally the centralized OPF cost for computing the optimality gap. YALMIP emphasizes modeling clarity and solver flexibility, and is suitable for rapid prototyping and debugging.
After receiving the user inputs, DistDCOPF.m loads all regional data, reconstructs inter-regional tie-lines, and augments each region’s branch matrix accordingly. Bus-angle variables, tie-line indices, and interface bus mappings are then initialized, along with the ADMM dual variables. During each ADMM iteration, every region solves its local DC OPF problem independently using the selected backend. The local objective minimizes the generation cost subject to regional constraints and tie-line consistency penalties. Following all local solves, the solver computes the angle mismatches on shared tie-lines, updates the dual variables, records the maximum primal residual, and checks the convergence condition. If a centralized optimal cost is available, the optimality gap is also computed and plotted. The procedure continues until convergence is reached or the iteration limit is met.
The solver supports any number of regions and enables reproducible distributed OPF studies. Its overall workflow is summarized in Algorithm 2.
Algorithm 2 DPLib Distributed OPF Solver
1: Input partitioned MATPOWER data, ADMM parameters rho, tolerance epsilon, maximum iterations
2: Load all regional data and global tie-line table
3: Augment each region with its corresponding inter-regional tie-line rows
4: Initialize regional variables, interface indices, and ADMM dual variables
5: Prepare regional consistency constraints and interface bus mappings
6:for each ADMM iteration do
7:  for each region do
8:   Solve regional OPF
9:  end for
10:  Compute tie-line angle mismatches across all region boundaries
11:  Update dual variables based on consistency violations
12:  Evaluate primal residual and check convergence criterion
13:  if convergence achieved then
14:   Break
15:  end if
16:end for
17: Visualize residuals, optimality gap (if applicable), and final regional solutions